OpenStudy (anonymous):

I am really stuck on this and need some help : Solve system : 4x1 + x2 = -5 16x1 + 4x2 = -20

OpenStudy (anonymous):

Are the numbers after x powers?

OpenStudy (anonymous):

Did you mean to put in y's for each second x as well?

OpenStudy (anonymous):

prime?

OpenStudy (lgbasallote):

are you just looking to solve for the system? because it will be no solution

OpenStudy (lgbasallote):

if that's algebra...

OpenStudy (anonymous):

no it says solve the system and its a matrix problem

OpenStudy (lgbasallote):

oh matrix lol :p

OpenStudy (anonymous):

the space for the answers says [x1] [x2] = [1] [s]

OpenStudy (anonymous):

linear algebra then?

OpenStudy (anonymous):

am I the only one noticing its the same equation twice... there are infinite solutions

OpenStudy (anonymous):

Do you guys know any calculus ? One of my answers was 0,1 for the bottom row and the answer was accepted

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No help att all? You guys say its simple I am giving the right format

OpenStudy (anonymous):

a possible answer is (x1,-5-4*x1), x1 (element of domain)

OpenStudy (anonymous):

What? I need the answer in a 2 x 2 matrix

OpenStudy (anonymous):

I know you probably did it right? But I do't get what that has to do with the answer?

OpenStudy (anonymous):

ok, so the matrix should represent the linear transformation?

OpenStudy (anonymous):

Yeah, I mean it says nothing about transformations but here is the exact wording Solve system 4x1 + x2 = -5 16x1 +4x2 = -20 [x1] = [x2] = 0 , 1 <------ solution I got for bottom row

OpenStudy (anonymous):

is something like this [4 1 -5] [16 4 -20]

OpenStudy (anonymous):

^YES Andresfon12

OpenStudy (anonymous):

Andresfron you got it!

OpenStudy (anonymous):

I had that too but

OpenStudy (anonymous):

I don't get how to solve by Gaussian for x1 and x2

OpenStudy (anonymous):

let think because did some problem like that last year

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

just row reduce the matrix

OpenStudy (anonymous):

If it helps I used Gaussian and got 1 , 1/4 <----- top row 0 , 0

OpenStudy (anonymous):

Right but after I row reduced its not accepting my answer how do I write it?

OpenStudy (amistre64):

what did you get for an answer? and what does your program accept for infinite solutions?

OpenStudy (anonymous):

I recall it

OpenStudy (anonymous):

Can you do out the reduction how do I work with the system to find x1 and x2 and write answer as 2 x 2 matrix

OpenStudy (anonymous):

need to multiply

OpenStudy (anonymous):

I got 1 1/4 0 0

OpenStudy (amistre64):

thats just the left side; do the right side as well

OpenStudy (anonymous):

No you need to row reduce the augmented matrix not just the matrix A

OpenStudy (anonymous):

ok right side is -5 60

OpenStudy (anonymous):

Then there is no solution.

OpenStudy (amistre64):

augment (A|b) into (I|x)

OpenStudy (amistre64):

ir something similar

OpenStudy (anonymous):

I am so lost

OpenStudy (anonymous):

row redudion at the bottom

OpenStudy (anonymous):

Can you write this with the numbers?

OpenStudy (amistre64):

you dont get 0+0 = 60 for starters

OpenStudy (anonymous):

Yeah I know my bottom row solution was 0, 1

OpenStudy (anonymous):

That's what the asker got.

OpenStudy (anonymous):

My bottom row solution of 0, 1 was accepted

OpenStudy (anonymous):

my top row was not accepted as correct thats what I mainly need help with

OpenStudy (anonymous):

i get like [0 4 0] for the bottom

OpenStudy (amistre64):

4x1 + x2 = -5 16x1 + 4x2 = -20 1 1/4 | -5/4 0 0 | 0

OpenStudy (anonymous):

^ I got that part Amistre

OpenStudy (anonymous):

sorry im wrong

OpenStudy (anonymous):

Here is a tip. The first row is a multiple of the second row. From this you can tell that there will be a zero row which means that there will be a free variable which mean.s that if there is a solution then there will be an infinite number of solutions

OpenStudy (amistre64):

\[\binom{x_1}{x^2}=\binom{-5/4}{0}+x_2\binom{1/4}{1}\]

OpenStudy (amistre64):

forgot my negative' -1/4

OpenStudy (anonymous):

i agree with amistre64

OpenStudy (amistre64):

one answer is [-5/4 , 0]

OpenStudy (anonymous):

Yes that was accepted!

OpenStudy (amistre64):

all the other answers are that + multiples of the other one

OpenStudy (anonymous):

i also agree with @amistre64

OpenStudy (anonymous):

but there is an infinite number of solutions.

OpenStudy (anonymous):

Thank you so much amistre

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Where was I going wrong?

OpenStudy (anonymous):

is all about multiply and dividing

OpenStudy (anonymous):

Like I know you probably have to go I am sorry

OpenStudy (amistre64):

without knowing what your program looks like, i wouldnt know where to begin; prolly tho in the parametric setup

OpenStudy (amistre64):

since we have a free variable; that portion is only equal to itself and nothing else ...

OpenStudy (anonymous):

Yeah that is true

OpenStudy (anonymous):

So there is a unique solution? only one solution?

OpenStudy (amistre64):

there is a .... homogenous solution and a particular solution , maybe lol

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

JAJAJA

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

I always thought rows of 0's means inconsistent but whatever i guess

OpenStudy (amistre64):

when does 0 equal 0 ?

OpenStudy (anonymous):

Wait I don't get it. What's the answer?

OpenStudy (anonymous):

read the defintion in your text book

OpenStudy (amistre64):

0=0 is always consistent 1=0 has one solution 0=1 never

OpenStudy (anonymous):

alright I will go over the concepts again

OpenStudy (anonymous):

Yeah you are right

OpenStudy (anonymous):

Gotcha bro.

OpenStudy (lgbasallote):

i'd love to say the answer is 42 but it's already been answered and it's too cliched =))